### v3

# Instantaneous Polarimetry with Zak-OTFS

Abstract
Polarimetry, which is the ability to measure the scattering response of the environment across orthogonal polarizations, is fundamental to enhancing wireless communication and radar system performance. In this paper, we utilize the Zak-OTFS modulation to enable instantaneous polarimetry within a single transmission frame. We transmit a Zak-OTFS carrier waveform and a spread carrier waveform mutually unbiased to it simultaneously over orthogonal polarizations. The mutual unbiasedness of the two waveforms enables the receiver to estimate the full polarimetric response of the scattering environment from a single received frame. Unlike existing methods for instantaneous polarimetry with computational complexity quadratic in the time-bandwidth product, the proposed method enables instantaneous polarimetry at near-linear complexity in the time-bandwidth product. Via numerical simulations, we show ideal polarimetric target detection and parameter estimation results with the proposed method, with improvements in computational complexity and greater clutter resilience over comparable baselines.
I Introduction
Polarimetry is an important tool for enhancing the performance of both wireless communication and radar systems. In wireless communication, polarimetry provides a diversity gain [1, 2, 3, 4], thereby improving the reliability of communication, as well as a spatial multiplexing gain [5, 6, 7, 8], which increases the capacity of the wireless link. Similarly, polarimetry increases the waveform degrees-of-freedom in radar systems [9, 10, 11, 12, 13, 14, 15, 16, 17, 18], providing more information about the target and enabling improved detection of targets with small radar cross section (RCS), such as drones.
Polarimetry is enabled in radar and communication systems by transmitting and receiving on two orthogonal polarizations, e.g., on vertical and horizontal polarizations. The receiver estimates the $2\times 2$ polarimetric scattering response of the wireless/radar channel across all four combinations of transmit and receive polarizations. A standard approach is to transmit polarized waveforms sequentially across two frames [11, 12, 13, 14, 15]; see Fig. 1(1(a)) for an example with frequency modulated continuous wave (FMCW) transmissions. From its measurements in each frame, the receiver estimates $2\times 1$ slices of the full $2\times 2$ polarimetric scattering response. Such an approach does not provide instantaneous estimates of the scattering response within a single frame. Changes in the scattering environment between the two frames (due to mobility) may partially decorrelate the obtained estimates [12, 13, 14, 15, 16]. Sequential polarimetry also prevents frame-by-frame processing & increases the system latency, which is a critical factor for radar and communication performance in highly dynamic environments. When utilizing continuous waveforms, such as FMCW and pulsed waveforms [19, 20, 21, 22], the computational complexity of sequential polarimetry is quadratic in the time-bandwidth product [23, 24, 25].
(a) Sequential Polarimetry with FMCW
(b) Instantaneous Polarimetry with Zak-OTFS (Ours)
Figure 1: Comparison of different approaches for polarimetry. (a) Sequential polarimetry with FMCW transmits polarized FMCW waveforms over two frames, with each frame subdivided into two halves with an up-chirp and a down-chirp respectively. The associated Doppler resolution is $\nicefrac{{2}}{{T}}$ and the computational complexity is $\mathcal{O}(B^{2}T^{2})$. (b) Instantaneous polarimetry with Zak-OTFS transmits a Zak-OTFS pulsone and a mutually unbiased spread waveform obtained via a unitary transformation of the pulsone in a single frame. Compared to the sequential approach in (a), the proposed approach has $2\times$ smaller latency, $2\times$ improved Doppler resolution of $\nicefrac{{1}}{{T}}$, and a computational complexity of only $\mathcal{O}(BT\log T)$.
To unlock the full benefits of polarimetry, it is crucial to estimate the $2\times 2$ polarimetric scattering response instantaneously within a single transmission frame. Previous work [13, 14, 15] has proposed transmitting mutually unbiased waveforms [^1] The term “mutually unbiased” is from quantum information theory [26]. Formally, two $d$-length waveforms are mutually unbiased if their inner product has magnitude $\nicefrac{{1}}{{\sqrt{d}}}$. Measurements from one waveform are “statistically independent” to those from the other waveform with uniform probability $\nicefrac{{1}}{{d}}$., i.e., waveforms with small inner products, simultaneously across orthogonal polarizations. Mutual unbiasedness ensures that the contribution of the other waveform looks like noise to the receiver when it projects its measurements onto the basis of one of the transmit waveforms. Projecting its measurements onto the basis of each of the two transmit waveforms provides the receiver with an estimate of different $2\times 1$ slices of the polarimetric scattering response; thus enabling full $2\times 2$ polarimetric scattering response estimation from a single received frame. Mutually unbiased waveforms have been designed in prior work [13, 14, 15] via phase-coding, i.e., by modulating a common carrier waveform, e.g., a rectangular waveform, with mutually unbiased sequences, e.g., Zadoff-Chu sequences with distinct roots [13] or complementary Golay pairs [14, 15]. While this approach offers excellent polarimetric target detection and parameter estimation (detailed later in Section IV), the computational complexity of polarimetry via phase-coding remains quadratic in the time-bandwidth product [25, 27]. Moreover, we have shown in [25] that separate selection of sequences and carrier waveforms, as done in phase-coding, may be sub-optimal from a radar waveform design perspective.
TABLE I: Comparison of different approaches for polarimetry; $B$ denotes signaling bandwidth and $T$ denotes frame interval.
Approach Frame(s) Target(s) Doppler Res. Complexity
Zak-OTFS (Ours) 1 $\mathbf{>1}$ $\mathbf{\nicefrac{{1}}{{T}}}$ $\mathbf{\mathcal{O}(BT\log T)}$
Phase-coded [13, 14, 15] 11 >1>1 $\nicefrac{{1}}{{T}}$ $\mathcal{O}(B^{2}T^{2})$
FMCW [12, 19] 22 11 $\nicefrac{{2}}{{T}}$ $\mathcal{O}(B^{2}T^{2})$
In this paper, we take an alternate approach to constructing mutually unbiased waveforms for instantaneous polarimetry by utilizing the Zak-OTFS (orthogonal time frequency space) modulation [28, 29, 30] in place of phase-coding. The carrier waveform in Zak-OTFS, termed pulsone, is a pulse localized in the delay-Doppler domain. We utilize a generalized discrete affine Fourier transform [25, 31] to transform the pulsone into a spread waveform that is mutually unbiased to the pulsone. Our approach has two advantages over phase-coding. First, scattering response estimation with pulsones and their unitary transformations is possible with near-linear computational complexity in the time-bandwidth product [25]. Second, unlike phase-coding, Zak-OTFS enables joint optimization of sequences and carrier waveforms, which has been shown to be optimal from a radar waveform design perspective in [25]. We illustrate our proposed approach in Fig. 1(1(b)) and describe it in more detail in Section III.
Table I places our contributions in the context of prior work, which is described in more detail in Section II. In addition to the advantages in computational complexity and latency, our proposed approach also improves upon the number of simultaneously detectable targets and the Doppler resolution over the sequential approach based on FMCW. Moreover, in Section IV we demonstrate that the proposed approach achieves ideal polarimetric target detection and parameter estimation [^2] The results in this paper serve as a proof-of-concept numerical demonstration of the concept; hardware demonstration will be pursued in future work. Potential challenges include managing the peak-to-average power ratio (PAPR) and over-the-air synchronization. With regards to the former, we note that a natural advantage of the proposed framework is its low PAPR of only $5.6$ dB; see [25, Section 4.2] and [31] for details. This reduction in PAPR has also been experimentally demonstrated in [32] in the sub-terahertz band. and greater clutter resilience compared to the phase-coded approach at smaller computational complexity.
Notation: $x$ denotes a complex scalar, $\mathbf{x}$ denotes a vector with $n$th entry $\mathbf{x}[n]$, and $\mathbf{X}$ denotes a matrix with $(n,m)$th entry $\mathbf{X}[n,m]$. $(\cdot)^{\ast}$ denotes complex conjugate, $(\cdot)^{\top}$ denotes transpose, $(\cdot)^{\mathsf{H}}$ denotes complex conjugate transpose and $\langle\mathbf{x},\mathbf{y}\rangle=\sum_{n}\mathbf{x}[n]\mathbf{y}^{\ast}[n]$ denotes the inner product. Calligraphic font $\mathcal{X}$ denotes operators or sets, with usage clear from context. $\emptyset$ denotes the empty set. $\mathbb{Z}$ denotes the set of integers and $\mathbb{Z}_{N}$ the set of integers modulo $N$. $(a,b)$ denotes the greatest common divisor of two integers $a,b$. $(\cdot)_{{}_{N}}$ denotes the value modulo $N$ and $(\cdot)^{-1}_{{}_{N}}$ denotes the inverse modulo $N$. $\delta(\cdot)$ denotes the delta function, $\delta[\cdot]$ denotes the Kronecker delta function and $\mathbf{I}_{N}$ denotes the $N\times N$ identity matrix.
II Polarimetry: Preliminaries
As described in the Introduction, polarimetry is enabled in radar and communication systems by transmitting and receiving on orthogonal polarizations, e.g., on vertical and horizontal polarizations, using dual-polarized antennas. Let $\mathsf{V}$ and $\mathsf{H}$ respectively denote vertical and horizontal polarization. We now describe how to model the polarimetric scattering response of a $P$-path wireless/radar channel. In uni-polarized systems, the channel gain of each path $p\in\{1,\cdots,P\}$ is modeled by a complex scalar $h^{(p)}$. With dual-polarized transmit and receive antennas, the channel gain is modeled instead by a $2\times 2$ polarimetric scattering response [10, 11, 12, 13, 14, 15]:
$\mathbf{H}^{(p)}$ $=\begin{bmatrix}h^{(p)}_{\mathsf{HH}}&h^{(p)}_{\mathsf{HV}}\\ h^{(p)}_{\mathsf{VH}}&h^{(p)}_{\mathsf{VV}}\end{bmatrix}=\mathbf{C}_{\mathsf{RX}}\mathbf{\Sigma}^{(p)}\mathbf{C}_{\mathsf{TX}},$ (1)
[…]
II-A Sequential Polarimetry via FMCW
[…]
In the specific case of sequential polarimetry via FMCW, it is known from [23, 24, 25] that cross-correlation-based channel estimation has computational complexity $\mathcal{O}(B^{2}T^{2})$, quadratic in the time-bandwidth product $BT$. Localizing a single path/target with FMCW requires subdividing each frame interval $T_{1}$ or $T_{2}$ into two halves and transmitting a chirp with positive slope (“up-chirp”) and a chirp with negative slope (“down-chirp”) in each half [23, 24], yielding an effective Doppler resolution of $\nicefrac{{2}}{{T}}$. Localizing multiple paths/targets with this approach [^4] It is possible to localize multiple paths/targets by subdividing each frame into four quarters with further degraded Doppler resolution of $\nicefrac{{4}}{{T}}$ [23, 24]. results in multiple false (“ghost”) targets [23, 24], degrading the detection performance. Moreover, the high sidelobes of FMCW makes detecting weak targets in the presence of stronger ones challenging [24, 25].
There are two primary drawbacks of the sequential approach independent of the drawbacks due to the choice of waveform. First, it requires the scattering environment to remain constant across two frames. Second, it prevents frame-by-frame processing. We now describe an approach for instantaneous polarimetry that overcomes these two drawbacks.
II-B Instantaneous Polarimetry via Phase-Coding
Instantaneous polarimetry overcomes the two drawbacks of sequential polarimetry by transmitting mutually unbiased waveforms simultaneously across both polarizations. Fig. 1(1(b)) illustrates the main idea. In a single frame interval, unit-norm waveforms $\mathbf{x}_{\mathsf{V}}(t)$ and $\mathbf{x}_{\mathsf{H}}(t)$ satisfying the property $\big|\int\mathbf{x}_{H}(t)\mathbf{x}_{V}^{*}(t-\tau)e^{-j2\pi\nu(t-\tau)}dt\big|\ll 1$ are transmitted in vertical and horizontal polarizations. The receiver estimates all four components of the $2\times 2$ matrix $\mathbf{H}^{(p)}$ from (1) by cross-correlating its dual-polarized received signals with the corresponding transmitted waveform. For example, the receiver estimates $h^{(p)}_{\mathsf{VH}}$ by cross-correlating its received signal in vertical polarization with $\mathbf{x}_{\mathsf{H}}(t)$.
Mutually unbiased waveforms have been designed in previous work [13, 14, 15] by phase coding a common carrier waveform with mutually unbiased discrete sequences, e.g., Zadoff-Chu sequences with distinct roots. The primary drawback of phase-coding is that the computational complexity of cross-correlation remains quadratic in the time-bandwidth product, $\mathcal{O}(B^{2}T^{2})$ [25, 27]. In the next Section, we describe how to design mutually unbiased waveforms using the Zak-OTFS modulation to enable instantaneous polarimetry with only $\mathcal{O}(BT\log T)$ computational complexity.
III Instantaneous Polarimetry via Zak-OTFS
We provide a brief overview of Zak-OTFS in the standard uni-polarized setting in Section III-A, referring the interested reader to [28, 29, 30] for a more detailed description of Zak-OTFS. We then extend the system model to polarimetry in Section III-B and detail our proposed approach in Section III-C.
III-A Overview of Zak-OTFS
The Zak-OTFS carrier waveform is a pulse in the delay-Doppler (DD) domain, formally a quasi-periodic localized function termed the DD pulsone [^5] Termed “pulsone” due to its structure of a pulse train modulated by a tone in the time domain, see [29, Fig. 2] for an illustration.. The DD pulsone is characterized by a delay period $\tau_{p}$ and a Doppler period $\nu_{p}$, with $\tau_{p}\nu_{p}=1$. The DD pulsone occupies infinite time and bandwidth. For practical implementation, the DD pulsone is limited to a time interval $T$ and a bandwidth $B$ via DD domain pulse shaping. The DD pulsone defines an orthonormal basis within the delay and Doppler periods with $BT=MN$ basis elements at $M=\nicefrac{{\tau_{p}}}{{\nicefrac{{1}}{{B}}}}=B\tau_{p}$ distinct locations along delay and $N=\nicefrac{{\nu_{p}}}{{\nicefrac{{1}}{{T}}}}=T\nu_{p}$ distinct locations along Doppler.
The DD pulsone is converted to a time domain (TD) waveform via the inverse Zak transform [28, 29, 30]. After Nyquist sampling, the discrete TD pulsone waveform is [28, 29, 30]:
$\mathbf{p}_{(k_{0},l_{0})}[n]$ $=\frac{1}{\sqrt{N}}\sum_{d\in\mathbb{Z}}e^{\frac{j2\pi}{N}dl_{0}}\delta[n-k_{0}-dM],$ (2)
where $k_{0}\in\mathbb{Z}_{M}$ indexes the location of the pulsone as a multiple of the delay resolution $\nicefrac{{1}}{{B}}=\nicefrac{{\tau_{p}}}{{M}}$, and $l_{0}\in\mathbb{Z}_{N}$ indexes the location of the pulsone as a multiple of the Doppler resolution $\nicefrac{{1}}{{T}}=\nicefrac{{\nu_{p}}}{{N}}$. The discrete TD signal on mounting $MN$ information symbols on the TD pulsones in (2) is:
$\mathbf{x}[n]$ $=\sum_{k_{0}=0}^{M-1}\sum_{l_{0}=0}^{N-1}\mathbf{X}[k_{0},l_{0}]\mathbf{p}_{(k_{0},l_{0})}[n],$ (3)
where $\mathbf{X}$ denotes the $M\times N$ array of information symbols.
After pulse shaping, the transmitted signal interacts with the scattering environment and is matched filtered at the receiver. The discrete TD received signal is given by [28, 31, 25]:
$\mathbf{y}[n]\!$ $=\!\sum_{k,l\in\mathbb{Z}_{MN}}\mathbf{h}_{\mathsf{eff}}[k,l]\mathbf{x}[(n-k)_{{}_{MN}}]e^{\frac{j2\pi}{MN}l(n-k)}\!+\!\mathbf{w}[n],$ (4)
where $\mathbf{h}_{\mathsf{eff}}[k,l]$ denotes the effective channel [^6] The effective channel approximates the physical channel when all paths are resolvable in delay with bandwidth $B$ and in Doppler with time $T$. that encompasses the effects of the physical scattering environment and transmit & receive pulse shaping/matched filtering [30, Eq. (7)], and $\mathbf{w}[n]$ denotes the additive noise at the receiver. For a scattering environment with $P$ paths/targets, let $\mathbf{h}_{\mathrm{phy}}(\tau,\nu)=\sum_{t=1}^{P}h^{(p)}\delta(\tau-\tau_{t})\delta(\nu-\nu_{t})$ denote the corresponding channel representation in the continuous DD domain. The effective channel is given by samples of the continuous effective channel, $\mathbf{h}_{\mathrm{eff}}[k,l]=\mathbf{h}_{\mathrm{eff}}\big(\tau=\frac{k\tau_{p}}{M},\nu=\frac{l\nu_{p}}{N}\big)$, where [28, 29, 30]:
$\mathbf{h}_{\mathrm{eff}}(\tau,\nu)$ $=\mathbf{w}_{{}_{\mathrm{RX}}}(\tau,\nu)*_{\sigma}\mathbf{h}_{\mathrm{phy}}(\tau,\nu)*_{\sigma}\mathbf{w}_{{}_{\mathrm{TX}}}(\tau,\nu).$ (5)
[…]
The effective channel is estimated at the receiver via the cross-ambiguity function [^8] When $\mathbf{y}=\mathbf{x}$, the expression $\mathbf{A}_{\mathbf{x},\mathbf{x}}[k,l]$ is called the self-ambiguity. [33, 24, 25, 31]:
$\widehat{\mathbf{h}}_{\mathsf{eff}}[k,l]$ $=\mathbf{A}_{\mathbf{y},\mathbf{x}}[k,l]$
$=\sum_{n=0}^{MN-1}\mathbf{y}[n]\mathbf{x}^{*}[(n-k)_{{}_{MN}}]e^{-\frac{j2\pi}{MN}l(n-k)},$ (6)
whose peaks in the absolute value indicate delay and Doppler bins of potential targets in the scattering environment. It has been shown in [24, 25] that computing the cross-ambiguity function in Zak-OTFS requires only $\mathcal{O}(BT\log T)$ complexity.
Accurate channel estimation is possible when the sequence $\mathbf{x}$ satisfies the crystallization condition [28, 29, 30, 33, 31]. Let $\mathcal{S}=\big\{(k,l)\big|\big|\mathbf{A}_{\mathbf{x},\mathbf{x}}[k,l]\big|=1\big\}$ denote the DD locations where the self-ambiguity function of $\mathbf{x}$ is unimodular, and let $\mathcal{C}$ denote the maximum DD support of the scattering environment999e.g., $\mathcal{C}=[k_{\min},k_{\max}]\times[l_{\min},l_{\max}]$ based on prior knowledge of the minimum/maximum delay and Doppler spreads of the scattering environment.. The crystallization condition requires:
$\bigg(\bigcup_{(k,l)\in\mathcal{S}}\big(\mathcal{C}+(k,l)\big)\bigg)\cap\bigg(\bigcup_{(k^{\prime},l^{\prime})\in\mathcal{S}}\big(\mathcal{C}+(k^{\prime},l^{\prime})\big)\bigg)=\emptyset,$ (7)
where $(k,l)\neq(k^{\prime},l^{\prime})$. In other words, translates of the channel support by locations where the self-ambiguity function is unimodular must not overlap for accurate channel estimation.
III-B Extension to Polarimetry
The system model in (4) is extended to polarimetry as:
$\mathbf{y}^{(j)}[n]\!$ $=\!\sum_{i\in\{\mathsf{V},\mathsf{H}\}}\!\sum_{k,l\in\mathbb{Z}_{MN}}\!\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]\mathbf{x}^{(i)}[(n-k)_{{}_{MN}}]e^{\frac{j2\pi}{MN}l(n-k)}$
$~~~~~~~~~~~~~~~~~~~~+\mathbf{w}^{(j)}[n],~i,j\in\{\mathsf{V},\mathsf{H}\},$ (8)
where $\mathbf{x}^{(i)}$ denotes the signal transmitted by the $i$-polarized transmit antenna, $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$ denotes the effective channel between the $i$-polarized transmit antenna and the $j$-polarized receive antenna, and $\mathbf{w}^{(j)}[n]$ denotes the additive noise at the $j$-polarized receive antenna. In (III-B), the polarimetric effective channel $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$ is defined similarly to (5) using the polarimetric continuous DD channel representation, $\mathbf{h}_{\mathrm{phy}}^{(j,i)}(\tau,\nu)=\sum_{t=1}^{P}h^{(p)}_{ji}\delta(\tau-\tau_{t})\delta(\nu-\nu_{t})$, where $h^{(p)}_{ji}$ is $(j,i)$th entry of the matrix $\mathbf{H}^{(p)}$ in (1).
III-C Proposed Approach for Instantaneous Polarimetry
We enable instantaneous polarimetry at near-linear complexity by designing mutually unbiased sequences via Zak-OTFS. To that end, in the following we define the generalized discrete affine Fourier transform (GDAFT) [25, 31], which maps pulsones in (2) to mutually unbiased spread waveforms. We have utilized the GDAFT to design radar waveform libraries in [25] and for spread carrier communication in [31].
Definition 1 ([25, 31])
The generalized discrete affine Fourier transform (GDAFT) of an $MN$-length sequence $\mathbf{x}$ is:
$\mathcal{F}_{a}\mathbf{x}[n]=\frac{1}{\sqrt{MN}}\sum_{m=0}^{MN-1}e^{\frac{j2\pi}{MN}(An^{2}+Bnm+Cm^{2})}\mathbf{x}[m],$
where $n\in\{0,\cdots,MN-1\}$, $A,B,C$ are co-prime to $MN$.
Theorem 1 ([25, 31])
The GDAFT in Definition 1 maps the discrete time pulsone in (2) localized at $(k_{0},l_{0})$ in the discrete DD domain to the spread carrier sequence:
$\mathbf{c}[n]=\mathcal{F}_{a}\mathbf{p}_{(k_{0},l_{0})}[n]=$ $\frac{e^{\frac{j2\pi}{MN}(An^{2}+Bnk_{0}+Ck_{0}^{2})}}{\sqrt{MN}}\epsilon_{N}\left(\frac{CM}{N}\right)_{J}$
$\times e^{-\frac{j2\pi}{N}(4CM)_{N}^{-1}(Bn+l_{0}+2Ck_{0})^{2}},$
where $\epsilon_{N}=1$ if $N\equiv 1\bmod 4$ & $\epsilon_{N}=j$ if $N\equiv 3\bmod 4$, and $\left(\frac{a}{b}\right)_{J}$ denotes the Jacobi symbol.
A useful consequence of Theorem 1 is that the output of the GDAFT is mutually unbiased to the pulsone [25, 31]:
$\mathbf{A}_{\mathbf{c},\mathbf{p}_{(k_{0},l_{0})}}[k,l]$ $=\frac{C_{(k_{0},l_{0})}[k,l]}{\sqrt{MN}},$ (9)
where $C_{(k_{0},l_{0})}[k,l]$ is a complex phase, $\big|C_{(k_{0},l_{0})}[k,l]\big|=1$, and $\mathbf{A}_{\mathbf{y},\mathbf{x}}[k,l]$ denotes the cross-ambiguity function as in (III-A). Moreover, the GDAFT preserves the near-linear computational complexity of cross-ambiguity-based channel estimation [25].
For instantaneous polarimetry, we transmit the pulsone and the output of the GDAFT in orthogonal polarizations, e.g.,
$\mathbf{x}^{(\mathsf{H})}[n]=\mathbf{p}_{(k_{0},l_{0})}[n],~\mathbf{x}^{(\mathsf{V})}[n]=\mathbf{c}[n].$ (10)
For accurate channel estimation, the GDAFT parameters $A,B,C$ in Theorem 1 are chosen such that $\mathbf{c}[n]$ satisfies the crystallization condition in (7) for all four components of the polarimetric effective channel $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$, for all $i,j\in\{\mathsf{V},\mathsf{H}\}$. For a detailed discussion on GDAFT parameter selection, see Fig. 1 and the associated example in [31, Section IV-C].
On obtaining the received signals per (III-B), the receiver computes the cross-ambiguity function per (III-A) between $\mathbf{y}^{(j)}$ and $\mathbf{x}^{(i)}$ to estimate the effective channel $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$, for all $i,j\in\{\mathsf{V},\mathsf{H}\}$. We now show how mutual unbiasedness per (9) enables accurate estimation of all four polarimetric effective channels $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$. Without loss of generality, we prove the result for the example considered in (10).
The estimate of $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$ from (III-A) and (III-B) is given by:
$\widehat{\mathbf{h}}_{\mathsf{eff}}^{(j,i)}[k,l]$ $=\mathbf{A}_{\mathbf{y}^{(j)},\mathbf{x}^{(i)}}[k,l]$
$=\sum_{n=0}^{MN-1}\mathbf{y}^{(j)}[n]\big(\mathbf{x}^{(i)}\big)^{*}[(n-k)_{{}_{MN}}]e^{-\frac{j2\pi}{MN}l(n-k)}$
$=\!\sum_{i^{\prime}\in\{\mathsf{V},\mathsf{H}\}}\!\sum_{k^{\prime},l^{\prime}}\mathbf{h}_{\mathsf{eff}}^{(j,i^{\prime})}[k^{\prime},l^{\prime}]\!\sum_{n=0}^{MN-1}\!\mathbf{x}^{(i^{\prime})}[(n-k^{\prime})_{{}_{MN}}]$
$~~~~e^{\frac{j2\pi}{MN}l^{\prime}(n-k^{\prime})}\big(\mathbf{x}^{(i)}\big)^{*}[(n-k)_{{}_{MN}}]e^{-\frac{j2\pi}{MN}l(n-k)}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\mathbf{A}_{\mathbf{w}^{(j)},\mathbf{x}^{(i)}}[k,l].$ (11)
For additive noise $\mathbf{w}^{(j)}$ uncorrelated with the transmitted signals $\mathbf{x}^{(i)}$, we have $\mathbf{A}_{\mathbf{w}^{(j)},\mathbf{x}^{(i)}}[k,l]=0$, for all $k,l$. On further making the substitution $n^{\prime}=(n-k^{\prime})_{{}_{MN}}$ we obtain:
$\widehat{\mathbf{h}}_{\mathsf{eff}}^{(j,i)}[k,l]$ $=\sum_{i^{\prime}\in\{\mathsf{V},\mathsf{H}\}}\!\sum_{k^{\prime},l^{\prime}}\mathbf{h}_{\mathsf{eff}}^{(j,i^{\prime})}[k^{\prime},l^{\prime}]\!\sum_{n^{\prime}=0}^{MN-1}\!\mathbf{x}^{(i^{\prime})}[n^{\prime}]e^{\frac{j2\pi}{MN}l^{\prime}n^{\prime}}$
$~~\big(\mathbf{x}^{(i)}\big)^{*}[(n^{\prime}-(k-k^{\prime})_{{}_{MN}})_{{}_{MN}}]e^{-\frac{j2\pi}{MN}l(n^{\prime}-(k-k^{\prime}))}$
$=\sum_{i^{\prime}\in\{\mathsf{V},\mathsf{H}\}}\!\sum_{k^{\prime},l^{\prime}}\mathbf{h}_{\mathsf{eff}}^{(j,i^{\prime})}[k^{\prime},l^{\prime}]e^{\frac{j2\pi}{MN}l^{\prime}(k-k^{\prime})}$
$~~~~~~~~~~\times\mathbf{A}_{\mathbf{x}^{(i^{\prime})},\mathbf{x}^{(i)}}[(k-k^{\prime})_{{}_{MN}},(l-l^{\prime})_{{}_{MN}}]$
$=\sum_{i^{\prime}\in\{\mathsf{V},\mathsf{H}\}}\mathbf{h}_{\mathsf{eff}}^{(j,i^{\prime})}[k,l]\ast_{\sigma}\mathbf{A}_{\mathbf{x}^{(i^{\prime})},\mathbf{x}^{(i)}}[k,l],$ (12)
where $\ast_{\sigma}$ denotes discrete twisted convolution [28, 29, 30, 33].
The expression in (III-C) is the sum of two terms:
$\widehat{\mathbf{h}}_{\mathsf{eff}}^{(j,i)}[k,l]$ $=\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]\ast_{\sigma}\mathbf{A}_{\mathbf{x}^{(i)},\mathbf{x}^{(i)}}[k,l]$
$+\mathbf{h}_{\mathsf{eff}}^{(j,\bar{i})}[k,l]\ast_{\sigma}\mathbf{A}_{\mathbf{x}^{(\bar{i})},\mathbf{x}^{(i)}}[k,l],$ (13)
where $\bar{i}$ denotes a polarization different from $i$ in the set $\{\mathsf{V},\mathsf{H}\}$. Since each sequence $\mathbf{x}^{(i)}$ satisfies the crystallization condition in (7), the first term is simply $\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]$. To simplify the second term, we substitute (9) to obtain:
$\widehat{\mathbf{h}}_{\mathsf{eff}}^{(j,i)}[k,l]$ $=\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l]+\mathbf{h}_{\mathsf{eff}}^{(j,\bar{i})}[k,l]\ast_{\sigma}\frac{C[k,l]}{\sqrt{MN}}$
$\approx\mathbf{h}_{\mathsf{eff}}^{(j,i)}[k,l],$ (14)
where $C[k,l]$ is a phase term similar to that in (9). Since the second term is the twisted convolution of the effective channel $\mathbf{h}_{\mathsf{eff}}^{(j,\bar{i})}[k,l]$ with a constant amplitude term, it simply raises the noise floor of the channel estimate. Computing each cross-ambiguity term only incurs $\mathcal{O}(BT\log T)$ complexity [^10] The complexity reduction from quadratic to near-linear is due to the symmetry of the Zak-OTFS carrier waveform. Specifically, the cross-ambiguity function computation with Zak-OTFS carrier waveforms reduces to an FFT (fast Fourier transform) calculation. For more details, see [25, Section 4.3]., and the overall complexity remains near-linear in $BT$.
IV Numerical Results
(a) FMCW, $\mathsf{HH}$
(b) Phase-coded, $\mathsf{HH}$
(c) Zak-OTFS, $\mathsf{HH}$
(d) FMCW, $\mathsf{VH}$
(e) Phase-coded, $\mathsf{VH}$
(f) Zak-OTFS, $\mathsf{VH}$
Figure 2: Heatmaps of estimated channels for a four-target environment with two targets with equal $h^{(p)}_{\mathsf{HH}}=0.7$ & $h^{(p)}_{\mathsf{HV}}=h^{(p)}_{\mathsf{VH}}=h^{(p)}_{\mathsf{VV}}=0$, and two targets with unequal $h^{(p)}_{\mathsf{HV}}=h^{(p)}_{\mathsf{VH}}\in\{0.3,0.95\}$ & $h^{(p)}_{\mathsf{HH}}=h^{(p)}_{\mathsf{VV}}=0$. (a) & (d): Sequential polarimetry via FMCW detects two false targets (“ghost targets”) in addition to the two true targets in the $\mathsf{HH}$ channel, and fails to detect the lower energy target in the $\mathsf{VH}$ channel due to the high sidelobes of the waveform. (b)-(c) & (e)-(f): Instantaneous polarimetry via mutually unbiased phase-coded and Zak-OTFS waveforms detects all four targets correctly in the $\mathsf{HH}$ and $\mathsf{VH}$ channels.
We now qualitatively and quantitatively compare the performance of the proposed approach from Section III with sequential polarimetry via FMCW (Section II-A) and instantaneous polarimetry via phase-coding (Section II-B). We also compare against uni-polarized systems. Our initial results are limited to target detection and estimation in the presence of noise; extensions to clutter are briefly pursued in Section IV-C.
[…]
IV-A Polarimetric Channel Estimation (Qualitative)
[…]
IV-B Target Detection & Parameter Estimation (Quantitative)
(a) FMCW, uni-polarized
(b) Phase-coded, uni-polarized
(c) Zak-OTFS, uni-polarized
(d) FMCW, dual-polarized
(e) Phase-coded, dual-polarized
(f) Zak-OTFS, dual-polarized
Figure 3: Histograms for single polarimetric target detection under the target present and target absent hypotheses. (a)-(c): Uni-polarization is insufficient for detecting polarimetric targets. (d): Dual-polarized FMCW is not optimal for polarimetric target detection due to high waveform sidelobes & false target detections (cf. Figs. 2(2(a)) & 2(2(d))). (e)-(f): Dual-polarized phase-coded and Zak-OTFS waveforms are optimal for polarimetric target detection (cf. Figs. 2(2(b))-(2(c)) & 2(2(e))-(2(f))).
(a) Detection ROC.
(b) Delay estimation RMSE.
(c) Doppler estimation RMSE.
Figure 4: Single target detection and estimation performance. (a): Receiver operating characteristic (ROC) curve showing ideal target detection with dual-polarized Zak-OTFS and phase-coded waveforms. The performance degrades with FMCW and/or uni-polarized waveforms. (b)-(c): Root mean squared error (RMSE) for delay and Doppler estimation, normalized by the corresponding delay and Doppler resolutions of $\nicefrac{{1}}{{B}}$ and $\nicefrac{{1}}{{T}}$. (b): Delay RMSE is similar for dual-polarized Zak-OTFS, phase-coded and FMCW systems at high signal-to-noise ratio (SNR), with significant improvements over uni-polarized waveforms. (c): Doppler RMSE is similar for dual-polarized Zak-OTFS and phase-coded waveforms at high SNR, with $\sim 1.5\times$ improvement over FMCW due to no loss in Doppler resolution. Significant improvements with dual-polarized vs uni-polarized waveforms.
We now quantify the performance of target detection and parameter (delay & Doppler) estimation. We consider a single target and model the entries of the $2\times 2$ matrix $\mathbf{H}^{(p)}$ as $h^{(p)}_{\mathsf{HH}}=a\sigma e^{j\phi}$, $h^{(p)}_{\mathsf{HV}}=h^{(p)}_{\mathsf{VH}}=a\sqrt{1-\sigma^{2}}e^{j\delta}$, $h^{(p)}_{\mathsf{VV}}=b\sigma e^{j\gamma}$, where $a,b\sim\mathsf{Bernoulli}(0.5)$ are i.i.d. symmetric Bernoulli random variables, $\sigma\sim U(0,1)$ is a standard uniform random variable, and $\phi,\delta,\gamma\sim U(0,2\pi)$ are i.i.d. uniform random variables in $0,2\pi)$. The delay & Doppler of the target are drawn uniformly at random according to $\tau_{t}\sim U(0,\nicefrac{{\tau_{p}}}{{4}})$, $\nu_{t}\sim U(-\nicefrac{{\nu_{p}}}{{8}},\nicefrac{{\nu_{p}}}{{8}})$. We generate $4\times 10^{4}$ Monte Carlo instances with signal-to-noise ratio (SNR) ranging from 20-20 dB to $20$ dB.
IV-B1 Target Detection
As the detection criteria, we compare the absolute value of the estimated channel at the delay-Doppler bin corresponding to the target’s location (representing the target present hypothesis) with the root-mean-squared value of the channel values at all locations other than the target’s location (representing the target absent hypothesis).
Fig. 3 plots the histograms corresponding to the two hypotheses for all considered systems. Overlapping histograms indicates poor target detectability. The histograms overlap significantly in the uni-polarized systems (Figs. 3(3(a))-(3(c))), since a single polarization is insufficient for estimating the full polarimetric scattering response. Polarimetry via FMCW also has significantly overlapping histograms (Fig. 3(3(d))) due to false detections and high sidelobes of the waveform. Polarimetry via phase-coding and Zak-OTFS has minimal overlap between the histograms (Figs. 3(3(e))-(3(f))), indicating their optimality.
Fig. 4(4(a)) plots the receiver operating characteristic (ROC) curve for all considered systems. Consistent with the findings from Fig. 3, we observe that polarimetry with phase-coded and Zak-OTFS waveforms achieves ideal target detection performance, with performance significantly degrading with FMCW and/or uni-polarized transmissions.
IV-B2 Parameter Estimation
For parameter estimation, we first detect peaks in the estimated channel following the procedure outlined in [33], which closely mimics the operation of a 2D constant false alarm rate detector from radar signal processing [21]. We threshold the energy of each channel location by the mean noise energy outside the region of interest $\mathsf{ROI}=[-\Delta\tau,\nicefrac{{\tau_{p}}}{{4}}+\Delta\tau]\times[-\nicefrac{{\nu_{p}}}{{8}}-\Delta\nu,\nicefrac{{\nu_{p}}}{{8}}+\Delta\nu]$ ($\Delta\tau$ and $\Delta\nu$ are guard widths to account for spread due to pulse shaping), scaled by an appropriate factor to achieve $10^{-6}$ false alarm rate [21]. After thresholding, the delay-Doppler location with the maximum channel energy is the estimated target location.
To fuse the target parameter estimates obtained across multiple polarimetric components, we compute their weighted average using entropy-based weights, $w^{(j,i)}=1-\nicefrac{{\mathcal{H}\big(\widehat{\mathbf{h}}_{\mathsf{eff}}^{(j,i)}\big)}}{{\log_{2}(MN)}}$ for all $i,j\in\{\mathsf{V},\mathsf{H}\}$, where $\mathcal{H}(\mathbf{h})=-\sum_{k,l}\big(\nicefrac{{|\mathbf{h}[k,l]|^{2}}}{{\sum_{k^{\prime},l^{\prime}}|\mathbf{h}[k^{\prime},l^{\prime}]|^{2}}}\big)\log_{2}\big(\nicefrac{{|\mathbf{h}[k,l]|^{2}}}{{\sum_{k^{\prime},l^{\prime}}|\mathbf{h}[k^{\prime},l^{\prime}]|^{2}}}\big)$ denotes the entropy [34] of a DD channel $\mathbf{h}$. Intuitively, such entropy-based weighting prioritizes parameter estimates obtained from polarimetric components with large variation in channel amplitudes (indicating the presence of targets) as opposed to those with little variation in channel amplitudes (indicating noise and the absence of any target). For illustration, consider the extreme case with no targets. Due to mutual unbiasedness, the estimated channel per (III-C) in this case has a constant energy level of $\nicefrac{{1}}{{MN}}$ and an entropy of $\log_{2}(MN)$, for which the chosen weight is $w^{(j,i)}=0$, i.e., such components are not allowed to bias our estimates.
Figs. 4(4(b))-(4(c)) plot the root-mean-squared error (RMSE) for delay and Doppler estimation for all considered systems. The RMSE for delay estimation is similar for polarimetry with FMCW, phase-coding and Zak-OTFS at high SNRs. The Doppler RMSE matches for polarimetry with phase-coding and Zak-OTFS, with $\sim 1.5\times$ improvement over FMCW, consistent with the explanation in Section II-A. Uni-polarized systems have significantly poorer delay and Doppler RMSE. Note that the RMSEs do not improve beyond a certain threshold due to the inherent resolution limits of the chosen waveforms. Future work will design optimal approaches achieving theoretical bounds, e.g., the Cramér-Rao bound [35].
IV-C Target Detection in Constant-$\gamma$ Clutter
Fig. 5 plots the detection ROC curve for single-target detection in constant-$\gamma$ clutter for all dual-polarized systems. For a metropolitan terrain and a carrier frequency of $f_{c}=4$ GHz, we obtain the clutter parameter value $\gamma=-1.99$ dB from [36]. Fig. 5 shows that instantaneous polarimetry via Zak-OTFS exhibits greater resilience to clutter as compared to methods based on phase-coded and FMCW transmissions.
Figure 5: Instantaneous polarimetry via Zak-OTFS exhibits greater resilience to clutter as compared to competing methods.
V Conclusion
In this paper, we proposed an alternate approach for instantaneous polarimetry using the Zak-OTFS modulation. We designed a spread carrier waveform mutually unbiased to the Zak-OTFS carrier waveform, and proposed to simultaneously transmit both waveforms over orthogonal polarizations. Unlike existing methods with computational complexity quadratic in the time-bandwidth product, the proposed method enables instantaneous polarimetry at near-linear complexity and greater clutter resilience. Future work will pursue experimental evaluation and applications to integrated sensing & communication.